OpenStudy (pagen13):

<1 and <2 are supplementary angles. m<1 is 4y+7 and m<2 is 9y+4. Find m<2

OpenStudy (anonymous):

1. 4y + 7 + 9y + 4 = 180 13y + 11 = 180 -11 -11 13y = 169 /13 /13 y = 13 9(13) + 4 117 + 4 = 121 m<2 = 121

OpenStudy (pagen13):

Thanks

OpenStudy (anonymous):

Please close post and if I helped please reward me with medal :)